# A logic of co-valuations

### v1

@@ 5.3. The Polish space $\rm{Mod}(L)$ of infinite compact $\omega$-structures in a countable signature $L$
Let $L$ be a countable signature. Fix contexts $X_{i}$, $i\in\mathbb{N}$, of size $|X_{0}|=1$, $|X_{i+1}|=2^{|X_{i}|}$, and let $\mathbb{P}=\bigsqcup X_{i}$. Let $M=(M,V)$ be a compact $\omega$-structure, and let $(v_{i})$ be a core of $V$. We can assume that $v_{i}\in V^{X_{i}}$. Let $a\in 2^{\mathbb{P}\times\mathbb{P}}$ be the characteristic function of the $\omega$-chain ordering on $\mathbb{P}$ given by $(v_{i})$. For $m\in\mathbb{N}$, let $T_{m}$ be the family of all the $m$-tuples in $X_{i}$, $i\in\mathbb{N}$. For $R\in L$, define $b_{R}\in 2^{T_{a(R)}}$ by
@@
Let $\alpha:\mathbb{N}\to C$ be an enumeration of the set $C$ defined by
@@
Observe that, by Remark 4.6 and Proposition 4.7, we can assume that there exists a function $\beta:\mathbb{N}\to\mathbb{N}$ that depends only on $L$, so that the following holds for $(v_{i})$. For every $m\in\mathbb{N}$, witnesses of Axiom (II) for elements represented by tuples $\alpha(0),\ldots,\alpha(m)$ can be found among consolidations of co-valuations $v_{0},\ldots,v_{\beta(m)}$. We associate with $\rm{Mod}(L)$ some fixed $\alpha$ and $\beta$ as above.
Let us consider $\mu=(a,b)\in 2^{\mathbb{P}\times\mathbb{P}}\times\prod_{R\in L}2^{T_{a(R)}}$ satisfying the following conditions:
@@
    $a$ is the characteristic function of an ordering relation $\leq_{\mu}$ on $\mathbb{P}$ turining it into an $\omega$-chain $(X_{i},\geq_{\mu}\upharpoonright X_{i}\times X_{i+1})$ such that each $\geq_{\mu}\upharpoonright X_{i+1}$ is co-bijective,
@@
      for all $j^{\prime}\geq j$, there is $(y_{k})\in X^{a(R)}_{j^{\prime}}$ such that $b_{R}((y_{k}))=1$, and, for $(x^{\prime}_{k})\in X^{a(R)}_{j}$, $x^{\prime}_{k}\geq_{\mu}y_{k}$ implies $x^{\prime}_{k}\in A_{k}$, $k\leq a(R)$.
Define co-valuations $v_{i}\in{\rm{CoV}}(X_{i},M)$, $i\in\mathbb{N}$, on $M={\mathsf{S}}(\mathbb{P},\leq_{\mu})$ by
@@
$R^{M}((S_{k}))\iff\forall(x_{k})\in T_{a(R)}\,((\forall k\,x_{k}\in S_{k})\rightarrow b_{R}((x_{k}))=1).$
Finally, put $V=\langle\{v_{i}\}\rangle$, and $M_{\mu}=(M,V)$. Observe that all $R^{M}$ are closed in $\tau_{V}$, so, by Corollary 4.14, $M_{\mu}$ is a compact $\omega$-structure.
Let us define $\rm{Mod}(L)$ as the subspace of $2^{\mathbb{P}\times\mathbb{P}}\times\prod_{R\in L}2^{T_{a(R)}}$ consisting of all the elements that satisfy Conditions (1) and (2) above. It is straightforward to verify that they are closed, i.e., that $\rm{Mod}(L)$ is a compact (in particular, Polish) subspace of $2^{\mathbb{P}\times\mathbb{P}}\times\prod_{R\in L}2^{T^{a(R)}}$. To sum up, we get
@@
Let $L$ be a countable signature. Then $\rm{Mod}(L)$ is a zero-dimensional compact space, and for every $\omega$-structure $M$, there exists $\mu\in\rm{Mod}(L)$ such that $M_{\mu}$ is isomorphic with $M$.
We can also consider a finer Polish topology on $\rm{Mod}(L)$. First of all,
@@
The implication from right to left is obivous. We show the converse. For any given $i\in\mathbb{N}$, $R\in L$, and tuple $\bar{x}=(x_{k})$ in $X^{a(R)}_{i}$, and for $j$, $\mathcal{A}$, and $(A_{k})$ as in (3), define $\succeq\in{\rm{CoV}}(\mathcal{A},X_{j})$ by
$A\succeq x\iff x\in A,$
and put $v=C_{\succeq}(v_{j})$. By (3a), we have $x_{k}]_{v_{i}}\supseteq\bigcup_{a\in A_{k}}[a]_{v}$. Using (3b), as in Proposition 3.12, we construct $x^{j}_{k}\in X_{i+j}$, $j\in\mathbb{N}$, $k\leq a(R)$ such that, $x^{0}_{k}=x_{k}$, $b_{R}((x^{j}_{k}))=1$, and, for $x^{\prime}\in X_{j}$, $x^{\prime}\geq_{\mu}x^{j+1}_{k}$ implies $x^{\prime}=x^{j}_{k}$. Then, since, by (1), $\mathbb{P}$ is an $\omega$-chain, we get that, for $x^{\prime}\in X_{i}$, $x^{\prime}\geq_{\mu}x^{k}_{j+1}$ implies $x^{\prime}=x^{0}_{k}=x_{k}$. In particular, every minimal selector $S_{k}\subseteq\{x^{j}_{k}\}^{\leq_{\mu}}$ is in $x_{k}]_{\pi_{v_{i}}}$. By the definition of $M_{\mu}$, we have $R^{M}((S_{k}))$, so $M_{\mu}\vDash_{v_{i}}R(\bar{x})$.
By Proposition 5.5, the sets $i,\geq]$, and $i,R,\bar{x}]$, where $i\in\mathbb{N}$, $\geq$ is a binary relation on $\bigsqcup_{i^{\prime}\leq i}X_{i^{\prime}}$, $R\in L$, and $\bar{x}\in X^{a(R)}$, defined by
$i,\geq]=\{\mu\in\rm{Mod}(L):(x\geq_{\mu}y\leftrightarrow x\geq y),\,x,y\in\bigsqcup_{i^{\prime}\leq i}X_{i^{\prime}}\},$
@@
form a sub-basis for the topology on $\rm{Mod}(L)$ inherited from $2^{\mathbb{P}\times\mathbb{P}}\times\prod_{R\in L}2^{T_{a(R)}}$. By standard arguments as in [7, Theorem 11.4.1], this topology can be refined to a Polish topology given by sub-basic open sets
@@
Fix $M$ as in the statement of the corollary, and suppose that it realizes a non-isolated type $p(X)$. By Corollary 5.8, there is a compact model $N$ of $T$ that omits $F\cup\{p(X)\}$. But then $M$ is isomorphic to $N$, a contradiction. ∎
@@
A theory $T$ in a countable signature has an atomic model if and only if isolated types are dense in $S_{n}(T)$ for every $n\in\mathbb{N}$.
@@
The implication from left to right is obvious. The converse follows from Theorem 5.6 because then non-isolated minimal types form a meager subset of $S_{n}(T)$ for every $n\in\mathbb{N}$. ∎
@@
We say that a theory is $\omega$-categorical if it has only one compact $\omega$-model.
@@
For sets $A$, $B$, and a relation $R$ of arity $m$ defined both on $A$, and on $B$, we say that a relation $\sqsupset\subseteq A\times B$ is $R$-preserving if, for every $\bar{b}\in B^{m}$, we have that $R(\bar{b})$ and $\bar{a}\sqsupset\bar{b}$ implies $R(\bar{a})$. It is $R$-surjective if, for every $\bar{a}\in A^{m}$, $R(\bar{a})$ implies that there is $\bar{b}\in B^{m}$ such that $R(\bar{b})$, and $\bar{a}\sqsupset\bar{b}$. For graphs $(G,\sqcap)$, $(H,\sqcap)$, we say that $\sqsupset\subseteq G\times H$ is edge-witnessing if, for $^{\prime}\in G, ggg\sqcap g^{\prime}$ implies that there is $h\in H$ with $g,g^{\prime}\sqsupset h$.
@@
The role of the graph relation $\sqcap$ is to control possible $\omega$-posets induced by sequences in subcategories of $\mathbf{S}_{L}$, as discussed in [3]. In particular, the category $\mathbf{S}_{\emptyset}$ is the category $\mathbf{B}$ described in Section 3 of [3].
Let $\mathbf{K}\subseteq\mathbf{S}_{L}$ be a category. We write $\mathbf{K}^{A}_{B}$ for the collection of all morphisms in $\mathbf{K}$ from $B$ to $A$. We say that $\mathbf{K}$ has amalgamation if for any $A,B_{0},B_{1}\in\mathbf{K}$, and $\sqsupset_{0}\in\mathbf{K}^{A}_{B_{0}}$, $\sqsupset_{1}\in\mathbf{K}^{A}_{B_{1}}$, there is $D\in\mathbf{K}$, and $\mathrel{\vphantom{\sqsupset}\text{\vtop{\halign{#\cr$\sqsupset$\cr$-$\cr}}}}_{0}\in\mathbf{K}^{B_{0}}_{D}$, $\mathrel{\vphantom{\sqsupset}\text{\vtop{\halign{#\cr$\sqsupset$\cr$-$\cr}}}}_{1}\in\mathbf{K}^{B_{1}}_{D}$ such that $\sqsupset_{0}\circ\mathrel{\vphantom{\sqsupset}\text{\vtop{\halign{#\cr$\sqsupset$\cr$-$\cr}}}}_{0}=\sqsupset_{1}\circ\mathrel{\vphantom{\sqsupset}\text{\vtop{\halign{#\cr$\sqsupset$\cr$-$\cr}}}}_{1}$. A sequence $(C_{i},\sqsupset_{i})$ in $\mathbf{K}$ is co-initial if for every $C\in\mathbf{K}$ there is $i\in\mathbb{N}$ and $\sqsupset\in\mathbf{K}^{C}_{C_{i}}$. It is absorbing if for every $i\in\mathbb{N}$, $C\in\mathbf{K}$, and $\sqsupset\in\mathbf{K}^{C_{i}}_{C}$ there is $j\in\mathbb{N}$, and $\mathrel{\vphantom{\sqsupset}\text{\vtop{\halign{#\cr$\sqsupset$\cr$-$\cr}}}}\mathbf{K}^{C}_{C_{j}}$ such that $\sqsupset\circ\mathrel{\vphantom{\sqsupset}\text{\vtop{\halign{#\cr$\sqsupset$\cr$-$\cr}}}}=\sqsupset^{i}_{j}$. It is Fraïssé if it is co-initial, and absorbing. One can easily show (see, e.g., [3, Proposition 4.30]) that if $\mathbf{K}$ is directed and has amalgamation, then there exists a Fraïssé sequence in $\mathbf{K}$.
As before, every sequence $(\mathbb{K}_{i},\sqsupset_{i})$ in $\mathbf{K}$ induces an $\omega$-chain poset $\mathbb{K}=(\bigsqcup\mathbb{K}_{i},\sqsupset)$. We can consider the induced compact $\omega$-structure $({\mathsf{S}}\mathbb{K},V)$ in signature $L$, where $V$ is generated by co-valuations $v_{i}\in\operatorname{\mathbf{Cov}}(\mathbb{K}_{i},M)$ determined by $\mathbb{K}^{\in}_{i}$, $i\in\mathbb{N}$, and, for $R\in L$ of arity $m$, we define
@@
for all $i\in\mathbb{N}$, and $x_{k}$, $k\leq m$, such that $x_{k}\in S_{k}$. Clearly, all $R^{M}$ are closed in $\tau_{V}$. We call $(\mathbb{K}_{i},\sqsupset_{i})$ $L$-faithful if $S(v_{i})=\mathbb{K}_{i}$, $i\in\mathbb{N}$. If $(\mathbb{K}_{i},\sqsupset_{i})$ is a Fraïssé sequence, we call $({\mathsf{S}}\mathbb{K},V)$ a Fraïssé limit of $\mathbf{K}$.
Note that because edge-witnessing morphisms are assumed to be wide in $\mathbf{K}$, and, by Proposition 3.12, $p^{\in}\neq\emptyset$, for $p\in\mathbb{K}$, we have that $G(v_{i})=(\mathbb{K}_{i},\sqcap)$ for every Fraïssé sequence $(\mathbb{K}_{i},\sqsupset_{i})$. The following fact can be proved in the same way as Proposition 3.12.
@@
It is known (see [3, Proposition 4.32]) that any two Fraïssé sequences in a category $\mathbf{K}\subseteq\mathbf{S}_{\emptyset}$ have homeomorphic spectra. It turns out that more can be said about them in the context of compact $\omega$-structures. The next two theorems suggest that for a second-countable compact $T_{1}$ space $X$ equipped with closed relations in signature $L$ that is topologically isomorphic to a Fraïssé limit of a category $\mathbf{K}\subseteq\mathbf{S}_{L}$, the (unique up to isomorphism) $\omega$-structure induced by an $L$-faithful Fraïssé sequence in $\mathbf{K}$ can be thought of as a canonical $\omega$-structure associated with $X$.
@@
It is straightforward to construct a formula $\mbox{Structure}_{K}(X)$ that says that $S(v)$ is isomorphic to a given structure $K$.
@@
Let $(N,W)$ a compact $\omega$-model of the theory $T$ of the Fraïssé limit $(M,V)$ of $(\mathbb{K}_{i},\sqsupset_{i})$ that omits all the types $D_{n}$. Then co-valuations $v$ such that $S(v)\in\mathbf{K}$ are co-initial in $W$. Observe that the statement of Lemma 6.4 can be expressed as an infinite conjunction of sentences in the logic of co-valuations, i.e., it is implied by $T$. Now, one easily constructs a core $(w_{i})$ of $W$ such that $(S(w_{i}),\geq^{w_{i}}_{w_{i+1}})$ is isomorphic with $(\mathbb{K}_{i},\sqsupset_{i})$. By Remark 4.16, $(N,W)$ is isomorphic with $(M,V)$. Finally, we apply Corollary 5.10.
Let $\mathbf{I}\subseteq\mathbf{S}_{\emptyset}$ be the category of all linear graphs with monotone morphisms, i.e., morphisms $\sqsupset\subseteq G\times H$ such that $A^{\sqsupset}$ is connected, provided that $A$ is connected. Let $\mathbf{P}\subseteq\mathbf{S}_{\emptyset}$ be the category of all linear graphs with all morphisms. Let $\mathbf{F}\subseteq\mathbf{S}_{\emptyset}$ be the category of all fans, i.e., trees with exactly one vertex of degree $\geq 3$, called the root, and spoke-monotone morphisms, i.e., morphisms that are monotone on every maximal path with the root as an end-point. All these categories are directed and have amalgamation (see [3, Proposition 5.13],[4, Theorem 4.14], and [3, Proposition 5.29]), and the respective Fraïssé limits are homeomorphic to the arc, the pseudo-arc, and the Lelek fan. Thus, we get

### v2

@@ 5.3. The Polish space $\rm{Mod}(L)$ of infinite compact $\omega$-structures in a countable signature $L$
Let $L$ be a countable signature. Fix pairwise disjoint contexts $X_{i}$, $i\in\mathbb{N}$, of size $|X_{0}|=1$, $|X_{i+1}|=2^{|X_{i}|}$, and let $\mathbb{P}=\bigcup X_{i}$. Let $M=(M,V)$ be an infinite compact $\omega$-structure, and let $(v_{i})$ be a core of $V$. We can assume that $v_{i}\in V^{X_{i}}$. Let $a\in 2^{\mathbb{P}\times\mathbb{P}}$ be the characteristic function of the $\omega$-chain ordering $\geq$ on $\mathbb{P}$ determined by $(v_{i})$, i.e.
$x\geq y\iff x\geq^{v_{i}}_{v_{j}}y,$
where $x\in X_{i}$, $y\in X_{j}$, $i\leq j$. For a fixed $m\in\mathbb{N}$, let $T_{m}$ be the family of all the $m$-tuples in $X_{i}$, $i\in\mathbb{N}$. For $R\in L$, we define $b_{R}\in 2^{T_{a(R)}}$ by
@@
We fix an enumeration $\alpha:\mathbb{N}\to C$ of the set $C$ defined by
@@
Observe that, by Remark 4.8 and Proposition 4.9, there exists a function $\beta:\mathbb{N}\to\mathbb{N}$, so that the following holds. For every compact $\omega$-structure $(M,V)$ in signature $L$ there exists a core $(v_{i})$ of $V$ such that, for every $m\in\mathbb{N}$, witnesses of Axiom (II) for elements represented by tuples $\alpha(0),\ldots,\alpha(m)$ can be found among consolidations of co-valuations $v_{0},\ldots,v_{\beta(m)}$. We fix such $\beta$, and consider only codes of compact $\omega$-structures in signature $L$ that satisfy the above condition (this will be needed to secure an appropriate topological structure on the set of codes). To be more specific, we consider $\mu=(a,b)\in 2^{\mathbb{P}\times\mathbb{P}}\times\prod_{R\in L}2^{T_{a(R)}}$ satisfying the following conditions:
@@
    $a$ is the characteristic function of an ordering relation $\leq_{\mu}$ on $\mathbb{P}$ turning it into an $\omega$-chain $(X_{i},\geq_{\mu}\upharpoonright X_{i}\times X_{i+1})$ such that each $\geq_{\mu}\upharpoonright X_{i+1}$ is co-bijective,
@@
      for all $j^{\prime}\geq j$, there is $(y_{k})\in X^{a(R)}_{j^{\prime}}$ such that $b_{R}((y_{k}))=1$, and, for all $k\leq a(R)$ and $x^{\prime}\in X_{j}$, we have that $x^{\prime}\geq_{\mu}y_{k}$ implies $x^{\prime}\in A_{k}$.
We define co-valuations $v_{i}\in{\rm{CoV}}(X_{i},M)$, $i\in\mathbb{N}$, on $M={\mathsf{S}}(\mathbb{P},\leq_{\mu})$ by
@@
$R^{M}((S_{k}))\iff\forall(x_{k})\in T_{a(R)}\,((\forall k\leq a(R)\,x_{k}\in S_{k})\rightarrow b_{R}((x_{k}))=1).$
Finally, we put $V=\langle\{v_{i}\}\rangle$, and $M_{\mu}=(M,V)$. Observe that all $R^{M}$ are closed in $\tau_{V}$, so, by Corollary 4.16, $M_{\mu}$ is a compact $\omega$-structure (note that Condition (3) has not been used so far).
Let us define $\rm{Mod}(L)$ as the subspace of $2^{\mathbb{P}\times\mathbb{P}}\times\prod_{R\in L}2^{T_{a(R)}}$ consisting of all the elements that satisfy Conditions (1)-(3) above. It is immediate to verify that these conditions are closed, i.e., that $\rm{Mod}(L)$ is a compact (in particular: Polish) subspace of $2^{\mathbb{P}\times\mathbb{P}}\times\prod_{R\in L}2^{T^{a(R)}}$. To sum up, we get
@@
Let $L$ be a countable signature. Then $\rm{Mod}(L)$ is a zero-dimensional metrizable compact space, and for every compact $\omega$-structure $M$, there exists $\mu\in\rm{Mod}(L)$ such that $M_{\mu}$ is isomorphic with $M$.
We will also need a finer Polish topology on $\rm{Mod}(L)$. First of all, we have:
@@
The implication from right to left is obvious. We show the converse. For any given $i\in\mathbb{N}$, $R\in L$, and tuple $\bar{x}=(x_{k})\in X^{a(R)}_{i}$, and for $j$, $\mathcal{A}$, and $(A_{k})$ as in Condition (3) above, we define $\succeq\in{\rm{CoV}}(\mathcal{A},X_{j})$ by
$A\succeq x\iff x\in A.$
We put $v=C_{\succeq}(v_{j})$. By (3a), we have $x_{k}]_{v_{i}}\supseteq\bigcup_{a\in A_{k}}[a]_{v}$. Using (3b), we construct, exactly as in the proof of Proposition 3.13, $x^{j}_{k}\in X_{i+j}$, $j\in\mathbb{N}$, $k\leq a(R)$ such that, $x^{0}_{k}=x_{k}$, $b_{R}((x^{j}_{k}))=1$, and, for $x^{\prime}\in X_{j}$, $x^{\prime}\geq_{\mu}x^{j+1}_{k}$ implies $x^{\prime}=x^{j}_{k}$. Then, since, by Condition (1), $(\mathbb{P},\leq_{\mu})$ is an $\omega$-chain, we get that, for $x^{\prime}\in X_{i}$, $x^{\prime}\geq_{\mu}x^{k}_{j+1}$ implies $x^{\prime}=x^{0}_{k}=x_{k}$. In particular, every minimal selector $S_{k}\subseteq\{x^{j}_{k}\}^{\leq_{\mu}}$ is an element of $x_{k}]_{v_{i}}$. By the definition of $M_{\mu}$, we get $R^{M}((S_{k}))$, so $M_{\mu}\vDash_{v_{i}}R(\bar{x})$.
Proposition 5.6 implies that the sets $i,\geq]$, and $i,R,\bar{x}]$, where $i\in\mathbb{N}$, $\geq$ is a binary relation on $\bigcup_{i^{\prime}\leq i}X_{i^{\prime}}$, $R\in L$, and $\bar{x}\in X^{a(R)}$, defined by
$i,\geq]=\{\mu\in\rm{Mod}(L):(x\geq_{\mu}y\leftrightarrow x\geq y),\,x,y\in\bigcup_{i^{\prime}\leq i}X_{i^{\prime}}\},$
@@
form a sub-basis for the topology on $\rm{Mod}(L)$ inherited from $2^{\mathbb{P}\times\mathbb{P}}\times\prod_{R\in L}2^{T_{a(R)}}$. By standard arguments as in the proof of [7, Theorem 11.4.1], this topology can be refined to a Polish topology given by sub-basic open sets
@@
Fix $M$ as in the statement of the corollary, and suppose that it realizes a non-isolated type $p$. By Corollary 5.9, there is a compact model $N$ of $T$ that omits $F\cup\{p\}$. But then $M$ is isomorphic to $N$, a contradiction. ∎
@@
A theory $T$ in a countable signature has an atomic model iff isolated types are dense in $S_{n}(T)$ for every $n\in\mathbb{N}$.
@@
The implication from left to right is obvious. The converse follows from Theorem 5.7 because then non-isolated types form a meager subset of $S_{n}(T)$ for every $n\in\mathbb{N}$. ∎
@@
We say that a theory is $\omega$-categorical if it has a unique up to isomorphism compact $\omega$-model.
@@
For sets $A$, $B$, and a relation $R$ of arity $m$ defined both on $A$, and on $B$, we say that a relation $\sqsupset\subseteq A\times B$ is $R$-preserving if, for every $\bar{b}\in B^{m}$, we have that $R(\bar{b})$ and $\bar{a}\sqsupset\bar{b}$ implies $R(\bar{a})$. It is $R$-surjective if, for every $\bar{a}\in A^{m}$, $R(\bar{a})$ implies that there is $\bar{b}\in B^{m}$ such that $R(\bar{b})$, and $\bar{a}\sqsupset\bar{b}$. For graphs $(G,\sqcap)$, $(H,\sqcap)$, we say that $\sqsupset\subseteq G\times H$ is edge-witnessing if, for all $^{\prime}\in G, ggg\sqcap g^{\prime}$ implies that there is $h\in H$ with $g,g^{\prime}\sqsupset h$.
@@
The role of the graph relation $\sqcap$ is to control the structure of $\omega$-posets induced by sequences in subcategories of $\mathbf{S}_{L}$, as discussed in [3]. In particular, the category $\mathbf{S}_{\emptyset}$ is the category $\mathbf{B}$ described in Section 3 of [3].
Let $\mathbf{K}\subseteq\mathbf{S}_{L}$ be a category. We write $\mathbf{K}^{A}_{B}$ for the collection of all morphisms in $\mathbf{K}$ from $B$ to $A$. We say that $\mathbf{K}$ has amalgamation if for any $A,B_{0},B_{1}\in\mathbf{K}$, and $\sqsupset_{0}\in\mathbf{K}^{A}_{B_{0}}$, $\sqsupset_{1}\in\mathbf{K}^{A}_{B_{1}}$, there are $D\in\mathbf{K}$, and $\mathrel{\vphantom{\sqsupset}\text{\vtop{\halign{#\cr$\sqsupset$\cr$-$\cr}}}}_{0}\in\mathbf{K}^{B_{0}}_{D}$, $\mathrel{\vphantom{\sqsupset}\text{\vtop{\halign{#\cr$\sqsupset$\cr$-$\cr}}}}_{1}\in\mathbf{K}^{B_{1}}_{D}$ such that $\sqsupset_{0}\circ\mathrel{\vphantom{\sqsupset}\text{\vtop{\halign{#\cr$\sqsupset$\cr$-$\cr}}}}_{0}=\sqsupset_{1}\circ\mathrel{\vphantom{\sqsupset}\text{\vtop{\halign{#\cr$\sqsupset$\cr$-$\cr}}}}_{1}$. A sequence $(C_{i},\sqsupset_{i})$ in $\mathbf{K}$ is co-initial if for every $C\in\mathbf{K}$ there are $i\in\mathbb{N}$ and $\sqsupset\in\mathbf{K}^{C}_{C_{i}}$. It is absorbing if for every $i\in\mathbb{N}$, $C\in\mathbf{K}$, and $\sqsupset\in\mathbf{K}^{C_{i}}_{C}$ there are $j\in\mathbb{N}$, and $\mathrel{\vphantom{\sqsupset}\text{\vtop{\halign{#\cr$\sqsupset$\cr$-$\cr}}}}\mathbf{K}^{C}_{C_{j}}$ such that $\sqsupset\circ\mathrel{\vphantom{\sqsupset}\text{\vtop{\halign{#\cr$\sqsupset$\cr$-$\cr}}}}=\sqsupset^{i}_{j}$. It is Fraïssé if it is co-initial, and absorbing. One can easily show (see, e.g., [3, Proposition 4.30]) that $\mathbf{K}$ is directed and has amalgamation iff there exists a Fraïssé sequence in $\mathbf{K}$.
As before, every sequence $(\mathbb{K}_{i},\sqsupset_{i})$ in $\mathbf{K}$ induces an $\omega$-chain poset $\mathbb{K}=(\bigsqcup\mathbb{K}_{i},\sqsupset)$. We can consider the induced compact $\omega$-structure $({\mathsf{S}}\mathbb{K},V)$ in signature $L$, where $V$ is generated by co-valuations $v_{i}\in\operatorname{\mathbf{Cov}}(\mathbb{K}_{i},M)$ determined by covers $\mathbb{K}^{\in}_{i}$, $i\in\mathbb{N}$, and, for $R\in L$ of arity $m$, we define
@@
for all $i\in\mathbb{N}$, and $x_{k}$, $k\leq m$, such that $x_{k}\in S_{k}$. Clearly, all $R^{M}$ are closed in $\tau_{V}$. We call $(\mathbb{K}_{i},\sqsupset_{i})$ $L$-faithful if the structure $S(v_{i})=\mathbb{K}_{i}$, $i\in\mathbb{N}$. If $(\mathbb{K}_{i},\sqsupset_{i})$ is a Fraïssé sequence, we call $({\mathsf{S}}\mathbb{K},V)$ a Fraïssé limit of $\mathbf{K}$.
Note that because edge-witnessing morphisms are assumed to be wide in $\mathbf{K}$, and, by Proposition 3.13, $p^{\in}\neq\emptyset$, for $p\in\mathbb{K}$, we have that the graph $G(v_{i})$ is equal to the reduct of $\mathbb{K}_{i}$ to $\{\sqcap\}$ for every Fraïssé sequence $(\mathbb{K}_{i},\sqsupset_{i})$. The following fact can be proved in the same way as Proposition 3.13.
@@
It is known (see [3, Proposition 4.32]) that any two Fraïssé sequences in a category $\mathbf{K}\subseteq\mathbf{S}_{\emptyset}$ have homeomorphic spectra. It turns out that more can be said about them in the context of compact $\omega$-structures. The next two theorems indicate that for a second-countable compact $T_{1}$ space $X$ equipped with closed relations in signature $L$ that is topologically isomorphic to a Fraïssé limit of a category $\mathbf{K}\subseteq\mathbf{S}_{L}$, the (unique up to isomorphism) $\omega$-structure induced by an $L$-faithful Fraïssé sequence in $\mathbf{K}$ can be thought of as a canonical $\omega$-structure associated with $X$.
@@
It is straightforward to construct a formula $\mbox{Structure}_{K}(X)$ that says that $S(v)$, where $v\in{\rm{CoV}}(X,M)$, is isomorphic to a given structure $K$.
@@
Let $(N,W)$ a compact $\omega$-model of the theory $T$ of the Fraïssé limit $(M,V)$ of $(\mathbb{K}_{i},\sqsupset_{i})$ that omits all the types $D_{n}$. Then co-valuations $w\in W$ such that $S(w)\in\mathbf{K}$ are co-initial in $W$. We observe that the statement of Lemma 6.4 can be expressed as an infinite conjunction of sentences in the logic of co-valuations, i.e., it is implied by $T$. Now, one easily constructs a core $(w_{i})$ of $W$ such that $(S(w_{i}),\geq^{w_{i}}_{w_{i+1}})$ is isomorphic with $(\mathbb{K}_{i},\sqsupset_{i})$. By Remark 4.18, $(N,W)$ is isomorphic with $(M,V)$. Finally, we apply Corollary 5.11.
Let $\mathbf{I}\subseteq\mathbf{S}_{\emptyset}$ be the category of all linear graphs with monotone morphisms, i.e., morphisms $\sqsupset\subseteq G\times H$ such that $A^{\sqsupset}$ is connected, provided that $A$ is connected. Let $\mathbf{P}\subseteq\mathbf{S}_{\emptyset}$ be the category of all linear graphs with all morphisms. Let $\mathbf{F}\subseteq\mathbf{S}_{\emptyset}$ be the category of all fans, i.e., trees with exactly one vertex of degree $\geq 3$, called the root, and spoke-monotone morphisms, i.e., morphisms that are monotone on every maximal path with the root as an end-point. All these categories are directed and have amalgamation (see [3, Proposition 5.13],[4, Theorem 4.14], and [3, Proposition 5.29]), and their respective Fraïssé limits are homeomorphic to the arc, the pseudo-arc, and the Lelek fan. Thus, we get